OpenStudy (anonymous):

3x^2y- x^2y Would it be 3x^2y orr 4x^2y? I'm thinking 4

OpenStudy (turingtest):

collect like terms\[\large3x^2y- x^2y=(3-1)x^2y=2x^2y\]

OpenStudy (anonymous):

2x^2 y